# Negate or Embrace: On How Misalignment Shapes Multimodal Representation Learning

### v1

@@ C.1 Lemmas
Clarification C.2 (Fixed Graded Lexicographic Order). The graded lexicographic order over index subsets of semantic variables is fixed purely for notational clarity and convenience. This choice imposes no constraints on the latent structure and is adopted without loss of generality. Since the true permutation of latent variables is unknown, any consistent ordering can be arbitrarily chosen to index the subsets associated with $\theta$ and $\rho$. Importantly, selection ($\theta$) and perturbation ($\rho$) biases are applied indirectly at the level of text observations—not via direct interventions on the latent space. As a result, the ordering merely determines how each value of $\theta$ or $\rho$ maps to a subset of semantic indices in $\mathbb{I}_{\mathbf{s}}$. For instance, omitting color information in captioning excludes the corresponding semantic variable; which value of $\theta$ encodes this depends on the chosen order. Fixing a graded lexicographic order therefore provides a deterministic and reproducible indexing scheme, without affecting the model’s generality or expressiveness.
C.4 Proof of Corollary 4.2
@@
Appendix D Numerical Simulation Details
@@
  • In § D.1, we outline the experimental setup, including hyperparameters, model architectures, and the construction of downstream tasks.
  • In § D.2, we present additional experiments that further validate our theoretical findings.
  • In § D.3, we analyze downstream performance under various perturbation bias settings.
D.1 Detailed Experimental Setup
@@
In the independent semantic variables setting, we set $\Sigma_{\mathbf{s}}=I_{10}$, so that $\mathbf{s}$ follows a factorized standard Gaussian distribution, consistent with nonlinear ICA settings [26, 60]. In the dependent semantic variables setting, we sample $\Sigma_{\mathbf{s}}$ from a Wishart distribution with an identity scale matrix and degrees of freedom equal to $10$, introducing an unknown causal structure among semantic latent variables, as in prior causal representation learning works [66, 12, 76].
@@
We parameterize the generative functions $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ using randomly initialized 3-layer invertible MLPs, following prior work [12]. Invertibility is enforced by maintaining a condition number threshold of $31\mathrm{e}{-}$ 3 for each layer.
@@
D.2 Additional Identification Results
@@
D.3 Additional Downstream Results
@@
Appendix E Experiment Details on MPI3D-Complex Dataset
We provide additional details on the MPI3D-Complex dataset that are not fully covered in § 5.2. In § E.1, we comprehensively describe the experimental setup, including a dataset overview, the selection and perturbation bias configurations used to generate text, model architecture, and training parameters. In § E.2, we present additional results, including the use of a linear classifier for predicting latent factors and an ablation study on encoder dimensionality.
@@ Appendix E Experiment Details on MPI3D-Complex Dataset
E.1 Detailed Experimental Setup
@@
The MPI3D-Complex dataset [18] contains 460,800 real-world images of resolution $64\times 64\times 3$, captured using a camera in a controlled environment. The dataset spans all possible combinations of seven mutually independent, discrete latent factors, as detailed in Table 6. Representative image samples are shown in Figure 11.
@@
For each setting, the dataset is partitioned into training, evaluation, and test subsets in a fixed ratio of $44{,}720:23{,}040:23{,}040$. Across all configurations, we train the image and text encoders for $200{,}000$ steps using the training subset, averaging results over three random seeds. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{\tiny MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
@@
For text encoding, we tokenize text using the nltk.PunktTokenizer, following the procedure in [12]. Tokenized sequences are transformed into two-dimensional one-hot embeddings and processed using a convolutional neural network (CNN) with a variable number of layers, determined by the shape of the tokenized input. The output dimensionality of the CNN is configured to match that of the image encoder, ensuring compatibility in the shared representation space.
@@ E.2 Additional Results
Appendix F Experiment Details on Causal3DIdent Dataset
We provide additional details on the Causal3DIdent dataset that are not fully covered in § 5.3. In § F.1, we offer a comprehensive description of the experimental setup, including the image and text latent factors, the image generation process, the design of selection and perturbation bias settings for text generation, and training configurations. In § F.2, we present supplementary results, including analyses of the learned text representations and assessments of the linearity of the learned embeddings.
F.1 Detailed Experimental Setup
@@
Following prior work [80, 66, 12, 76], we utilize the Causal3DIdent dataset to synthesize images from a predefined latent causal structure. Images are generated using the Blender renderer [8], which applies a complex rendering function parameterized by 11 input variables. In our configuration, the object’s $z$-position is fixed, leaving 10 latent factors that govern image generation.
@@
To introduce text-specific variation, we employ five manually designed templates to generate text from the latent factors under each bias setting, adapting the text rendering pipeline from [12]. A complete list of latent factors and their types is provided in Table 9.
@@
Across all experimental settings, we train the image and text encoders for $100{,}000$ steps on the training subset, using three different random seeds to ensure robustness. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{\tiny MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
@@ F.2 Additional Results
Regarding the text-specific factor phrase, we find it to be partially encoded in the learned representations. This contrasts with the image-specific continuous factors, which are consistently omitted. The partial identifiability of phrase is likely attributable to its discrete nature, which violates the conditions typically required for modality-specific factors to be excluded—consistent with findings reported in [12]. Moreover, this effect appears more pronounced under selection bias settings, particularly when the encoding dimensionality exceeds the true dimensionality of the unbiased semantic subspace.
@@
Appendix G Computation Resources
@@ Appendix H Discussions
H.1 Modeling Randomly Missing Latent Semantic Variables
Our current framework addresses misalignment arising primarily from fixed selection and perturbation biases in textual annotations. However, real-world datasets—particularly large-scale, user-generated corpora such as LAION-5B [58]—often display random and unstructured semantic omissions. Extending our latent variable model to accommodate such random missingness (e.g., missing completely at random or missing not at random) poses both theoretical and practical challenges. Future research should examine the identifiability consequences of randomly missing semantic variables and establish conditions that ensure partial or probabilistic recovery of latent factors remains feasible.
H.2 Toward Linearity of Multimodal Representations
@@ H.3 Broader Implications
Our findings offer computational support for a contemporary perspective on linguistic relativity [68], which posits that language shapes human perception and conceptualization. In multimodal AI, linguistic supervision implicitly determines which aspects of the visual world are foregrounded—and which are omitted. Selection and perturbation biases in textual annotations thus act as epistemic filters, constraining the conceptual space that models can represent and generalize over. This reframes dataset design as both an epistemic and normative act: inclusion or omission in annotations encodes a stance on salience, relevance, or appropriateness. Two practical insights follow. First, achieving generalizable representations requires supervision that faithfully captures the intended semantic scope without systematic omissions. Second, curating data to intentionally include or exclude ethically sensitive or socially salient content provides a mechanism for aligning learned representations with human values [3, 10]. When annotation is treated as an epistemic commitment, dataset design becomes a central tool for shaping both semantic fidelity and ethical alignment.
@@
Beyond explicit annotation content, human annotation errors—such as consistent omissions or mislabels—reveal subtle but powerful behavioral signals. These errors are not uniformly random; rather, they concentrate in dimensions that humans intuitively deprioritize under limited attention [63, 23]. Core attributes like risk or intentionality are rarely mislabeled, while peripheral or low-salience details are more frequently neglected. Such patterns implicitly encode a value hierarchy over semantic factors. This perspective invites a reinterpretation: rather than discarding annotation errors as noise, we can study them as reflections of human value structure. Systematically neglected factors often carry little perceived cost when missed, suggesting lower social or cognitive importance. These regularities form a behavioral prior—informing models not only what remains invariant to environment changes, but what is valued. Incorporating such cues enables a richer form of alignment grounded in human preferences, priorities, and ethical sensibilities [16].

### v2

@@ B.1 Lemmas
Clarification B.2 (Fixed Graded Lexicographic Order).
The graded lexicographic order over subsets of semantic indices is fixed solely for notational consistency and clarity. For example, given semantic variables $\mathbf{s}=(s_{1},s_{2},s_{3})$ with index set $\mathbb{I}_{\mathbf{s}}=\{1,2,3\}$, the graded lexicographic order over the non-empty powerset $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$ yields: $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})=\{\{1\},\{2\},\{3\},\{1,2\},\{1,3\},\{2,3\},\{1,2,3\}\}.$ This ordering imposes no constraints on the underlying latent structure and is adopted without loss of generality. Since the true permutation of latent variables is unidentifiable, any consistent order can be used to index subset-valued parameters such as $\theta$ (selection) and $\rho$ (perturbation). Importantly, both $\theta$ and $\rho$ act indirectly through text observations rather than directly operating on the latent space. The fixed order simply determines how each value of $\theta$ or $\rho$ maps to a subset of semantic indices. For instance, omitting color in a caption corresponds to removing the associated latent variable—which index encodes this depends on the fixed ordering. Thus, adopting a graded lexicographic order ensures a deterministic and reproducible indexing scheme without limiting model generality or expressiveness.
B.4 Proof of Corollary 4.2
@@
Appendix C Numerical Simulation Details
@@
  • In § C.1, we outline the experimental setup, including hyperparameters, model architectures, and the construction of downstream tasks.
  • In § C.2, we present additional experiments that further validate our theoretical findings.
  • In § C.3, we analyze downstream performance under various perturbation bias settings.
C.1 Detailed Experimental Setup
@@
In the independent semantic variables setting, we set $\Sigma_{\mathbf{s}}=I_{10}$, so that $\mathbf{s}$ follows a factorized standard Gaussian distribution, consistent with nonlinear ICA settings (Khemakhem et al., 2020; Sorrenson et al., 2020). In the dependent semantic variables setting, we sample $\Sigma_{\mathbf{s}}$ from a Wishart distribution with an identity scale matrix and degrees of freedom equal to $10$, introducing an unknown causal structure among semantic latent variables, as in prior causal representation learning works (Von Kügelgen et al., 2021; Daunhawer et al., 2022; Yao et al., 2023).
@@
We parameterize the generative functions $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ using randomly initialized 3-layer invertible MLPs, following prior work (Daunhawer et al., 2022). Invertibility is enforced by maintaining a condition number threshold of $31\mathrm{e}{-}$ 3 for each layer.
@@
C.2 Additional Identification Results
@@
C.3 Additional Downstream Results
@@
Appendix D Experiment Details on MPI3D-Complex Dataset
We provide additional details on the MPI3D-Complex dataset that are not fully covered in § 5.2. In § D.1, we comprehensively describe the experimental setup, including a dataset overview, the selection and perturbation bias configurations used to generate text, model architecture, and training parameters. In § D.2, we present additional results, including the use of a linear classifier for predicting latent factors and an ablation study on encoder dimensionality.
@@ Appendix D Experiment Details on MPI3D-Complex Dataset
D.1 Detailed Experimental Setup
@@
The MPI3D-Complex dataset (Gondal et al., 2019) contains 460,800 real-world images of resolution $64\times 64\times 3$, captured using a camera in a controlled environment. The dataset spans all possible combinations of seven mutually independent, discrete latent factors, as detailed in Table 6. Representative image samples are shown in Figure 11.
@@
For each setting, the dataset is partitioned into training, evaluation, and test subsets in a fixed ratio of $44{,}720:23{,}040:23{,}040$. Across all configurations, we train the image and text encoders for $200{,}000$ steps using the training subset, averaging results over three random seeds. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
@@
For text encoding, we tokenize text using the nltk.PunktTokenizer, following the procedure in (Daunhawer et al., 2022). Tokenized sequences are transformed into two-dimensional one-hot embeddings and processed using a convolutional neural network (CNN) with a variable number of layers, determined by the shape of the tokenized input. The output dimensionality of the CNN is configured to match that of the image encoder, ensuring compatibility in the shared representation space.
@@ D.2 Additional Results
Appendix E Experiment Details on Causal3DIdent Dataset
We provide additional details on the Causal3DIdent dataset that are not fully covered in § 5.3. In § E.1, we offer a comprehensive description of the experimental setup, including the image and text latent factors, the image generation process, the design of selection and perturbation bias settings for text generation, and training configurations. In § E.2, we present supplementary results, including analyses of the learned text representations and assessments of the linearity of the learned embeddings.
E.1 Detailed Experimental Setup
@@
Following prior work (Zimmermann et al., 2021; Von Kügelgen et al., 2021; Daunhawer et al., 2022; Yao et al., 2023), we utilize the Causal3DIdent dataset to synthesize images from a predefined latent causal structure. Images are generated using the Blender renderer (Community, 2018), which applies a complex rendering function parameterized by 11 input variables. In our configuration, the object’s $z$-position is fixed, leaving 10 latent factors that govern image generation.
@@
To introduce text-specific variation, we employ five manually designed templates to generate text from the latent factors under each bias setting, adapting the text rendering pipeline from (Daunhawer et al., 2022). A complete list of latent factors and their types is provided in Table 9.
@@
Across all experimental settings, we train the image and text encoders for $100{,}000$ steps on the training subset, using three different random seeds to ensure robustness. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
@@ E.2 Additional Results
Regarding the text-specific factor phrase, we find it to be partially encoded in the learned representations. This contrasts with the image-specific continuous factors, which are consistently omitted. The partial identifiability of phrase is likely attributable to its discrete nature, which violates the conditions typically required for modality-specific factors to be excluded—consistent with findings reported in (Daunhawer et al., 2022). Moreover, this effect appears more pronounced under selection bias settings, particularly when the encoding dimensionality exceeds the true dimensionality of the unbiased semantic subspace.
@@
Appendix F Computation Resources
@@ Appendix G Discussions
G.1 Modeling Randomly Missing Latent Semantic Variables
Our current framework addresses misalignment arising primarily from fixed selection and perturbation biases in textual annotations. However, real-world datasets, particularly large-scale, user-generated corpora such as LAION-5B (Schuhmann et al., 2022), often display random and unstructured semantic omissions. Extending our latent variable model to accommodate such random missingness (e.g., missing completely at random or missing not at random) poses both theoretical and practical challenges. Future research should examine the identifiability consequences of randomly missing semantic variables and establish conditions that ensure partial or probabilistic recovery of latent factors remains feasible.
G.2 Toward Linearity of Multimodal Representations
@@ G.3 Broader Implications
Our findings offer computational support for a contemporary perspective on linguistic relativity (Whorf, 2012), which posits that language shapes human perception and conceptualization. In multimodal AI, linguistic supervision implicitly determines which aspects of the visual world are foregrounded—and which are omitted. Selection and perturbation biases in textual annotations thus act as epistemic filters, constraining the conceptual space that models can represent and generalize over. This reframes dataset design as both an epistemic and normative act: inclusion or omission in annotations encodes a stance on salience, relevance, or appropriateness. Two practical insights follow. First, achieving generalizable representations requires supervision that faithfully captures the intended semantic scope without systematic omissions. Second, curating data to intentionally include or exclude ethically sensitive or socially salient content provides a mechanism for aligning learned representations with human values (Bender and Friedman, 2018; Crawford, 2021). When annotation is treated as an epistemic commitment, dataset design becomes a central tool for shaping both semantic fidelity and ethical alignment.
@@
Beyond explicit annotation content, human annotation errors—such as consistent omissions or mislabels—reveal subtle but powerful behavioral signals. These errors are not uniformly random; rather, they concentrate in dimensions that humans intuitively deprioritize under limited attention (Tversky and Kahneman, 1990; Itti and Koch, 2001). Core attributes like risk or intentionality are rarely mislabeled, while peripheral or low-salience details are more frequently neglected. Such patterns implicitly encode a value hierarchy over semantic factors. This perspective invites a reinterpretation: rather than discarding annotation errors as noise, we can study them as reflections of human value structure. Systematically neglected factors often carry little perceived cost when missed, suggesting lower social or cognitive importance. These regularities form a behavioral prior—informing models not only what remains invariant to environment changes, but what is valued. Incorporating such cues enables a richer form of alignment grounded in human preferences, priorities, and ethical sensibilities (Gabriel, 2020).